# Towards Universal Dense Blocking for Entity Resolution

### v1

@@ 5.3. Comparison with Sparse Blocking Methods
Initially, we compare the effectiveness of UBlocker and SOTA sparse blocking methods. Our experimental results, shown in Table 4, reveal that Sparkly surpasses UBlocker by about 2% mAP on average. However, it’s worth noting that UBlocker is comparable to Sparkly on simple homogeneous datasets, and blocking better on some more difficult datasets, such as “walmart-amazon_heter” and “notebook”. In addition, the ensemble of UBlocker and Sparkly can improve PC by up to 5% on dataset “cora”. Lastly, self-supervised dense retrieval models have outperformed BM25 as more high quality data become available and model parameters expand. This suggests that the performance of UBlocker will continue to improve over time, showing great promise for universal self-supervised dense blocking.
Finding 4. UBlocker is comparable to SOTA sparse blocking methods and they complement each other to a certain extent.
@@
Furthermore, we perform efficiency analysis on scalability datasets from our proposed benchmark. To this end, we downsample the entity collections from $10^{2}$ to $10^{6}$ and compare the blocking time of different approaches. As shown in Figure 4, sparse blocking methods are efficient for small numbers of records, but the difference decreases as the number of records increases. We can see that UBlocker has the slowest time growth curve because it generates fixed-length dense representations for records. However, for Sparkly with a 3-gram tokenizer, the number of tokens and the number of record pairs with the same tokens grow non-linearly, posing challenges for index storage and nearest neighbor search. Therefore, q-gram technology improves the effectiveness but sacrifices some scalability at larger scales.
Finding 5. Sparse blocking methods are more efficient than UBlocker for small numbers of records, but the difference decreases as the number of records increases.
@@ 5.4. Ablation Study
We now perform ablation studies on UBlocker to validate the importance of specific components, as depicted in Table 5.
Table 5. Average performance in ablation studies. UBlocker “w/o OD” means pre-training with sufficient out-domain data instead of open-domain GitTables. “RE”, “LP”, “FP”, and “DD” stand for record embedding, literal paraphrasing, feature paraphrasing, and duplicate detection, respectively.
@@
UBlocker w/o OD 46.48 89.37 18.00 35.76
UBlocker w/o RE 48.31 89.29 19.72 33.82
UBlocker w/o LP 49.09 89.40 20.23 34.71
UBlocker w/o FP 48.64 89.49 19.75 35.94
UBlocker w/o DD 48.31 89.06 19.65 36.12
UBlocker 49.60 89.86 20.29 32.35
Effectiveness of Pre-training Data. To explore the effect of pre-training data diversity on UBlocker, we replace the training data extracted from open-domain GitTables with out-domain tables source from scalability datasets in our benchmark. This helps to establish whether the pre-training data serves solely to bridge the gap between natural language and structured records, or if it also contributes to the universality of UBlocker. Experimental results in Table 5 show that UBlocker achieves a significant performance gain over the model trained on out-domain data, suggesting that open-domain data pre-training is crucial for its universality.
@@
Effectiveness of Data Paraphrasing. We then experiment with the effect of different data paraphrasing approaches by removing them in turn. As shown in Table 5 again, we find that both feature paraphrasing (FP) and lexical paraphrasing (LP) contribute to the performance of UBlocker, among which LP is more pronounced because this type of paraphrasing is more refined, diverse, and stable, without the chance of introducing false positives. Moreover, due to the complexity of the real data, the application of duplicate detection (DD) achieves a 1.29% mAP improvement.
Finding 7. Pre-training on open-domain data and the techniques applied in UBlocker are effective and necessary.
@@ 7. Conclusion
The main focus of this paper is to explore the potential of developing a universal dense blocker that can be applied across diverse scenarios and domains through large-scale self-supervised learning. To achieve this, we propose a unified contrastive pre-training framework that exploits the power of cross-domain pre-training for developing the universal dense blocker, UBlocker. The pre-training is performed via self-supervised contrastive learning on large-scale relational tables. We also introduce a new record embedding technique that allows Transformer-based architectures to better model individual records. In addition, we propose three data paraphrasing techniques to generate positives for training. To assess the effectiveness and universality of UBlocker, we create a new benchmark for universal blocking evaluation that covers a wide range of blocking tasks from multiple domains and scenarios. Experimental results demonstrate the effectiveness and universality of UBlocker, holding great promise for universal dense blocking.
@@ Appendix B Benchmark Discussion
Our proposed benchmark value can be further validated by experimental results. On the one hand, Table 4 shows that existing blocking methods can effectively address the homogeneous and simple datasets that previous studies mainly focused on in Table 2. There is still room for improvement on some of our newly introduced datasets, providing new challenges and opportunities for future research. On the other hand, our proposed benchmark is the first to evaluate the general performance of different methods in different scenarios using the same parameters on all datasets, laying the foundation for future automatic, data-driven sparse blocking tuning approaches or good attribute identification for blocking. Finally, by comparing the performance differences of UBlocker on artificially and naturally heterogeneous datasets with their corresponding homogeneous datasets, we can see that the mainstream approach of artificially constructing dirty data does not pose much of a challenge to deep models. How to better perform entity resolution directly on real highly heterogeneous data (such as “walmart-amazon_heter”) is also a future research direction.
@@ 13             $candidate\_pair\leftarrow(j,indices[j][i])$
This section presents how to use UBlocker for blocking any number of entity record collections. It is worth mentioning that our general implementation of the nearest neighbor blocker can apply different sparse or dense vectorizers and indexers for a thorough and fair comparison of different methods. As shown in Algorithm 2, the nearest neighbor blocker operates on a list of entity collections, using a record vectorizer and a nearest neighbor indexer to identify $k$ nearest neighbors for each record. In cases with more than two collections, they are unified into a single one. The first and last collections are vectorized as the $source$ and $target$ respectively, with the latter being indexed. The algorithm iteratively searches for $k$ nearest neighbors of source vectors and stores these indices. Subsequently, candidate pairs are formulated, each constituted of a source index and its corresponding nearest neighbor index from the target. Each candidate pair is prioritized and deduplicated according to its nearest neighbor ranking, to facilitate progressive entity matching or evaluation. This sequence of candidate pair sets forms the final output of the algorithm.
Besides offering a more precise assessment for blocking, our suggested benchmark value can be further corroborated by experimental results. On one hand, Table 4 demonstrates that current sparse and dense blocking methods effectively tackle homogeneous and simple datasets primarily studied in Table 2. However, there is scope for improvement on some newly introduced datasets, which present new challenges and opportunities for future research. On the other hand, our suggested benchmark is the first to evaluate general performance across various scenarios using the same parameters for all datasets, setting the groundwork for future data-driven sparse blocking tuning approaches (Papadakis et al., 2022) or more versatile dense blocking model evaluations. Lastly, by comparing UBlocker’s performance differences on artificially and naturally heterogeneous datasets with their homogeneous counterparts, we observe that artificial dirty data construction does not significantly challenge deep models. A future research direction includes improving entity resolution on highly heterogeneous real data (such as ”walmart-amazon_heter”).

### v2

@@ 5.3. Comparison with Sparse Blocking Methods
Initially, we compare the effectiveness of UniBlocker and SOTA sparse blocking methods. Our experimental results, shown in Table 4, reveal that Sparkly surpasses UniBlocker by about 2% mAP on average. However, it’s worth noting that UniBlocker is comparable to Sparkly on simple homogeneous datasets, and blocking better on some more difficult datasets, such as “walmart-amazon_heter” and “notebook”. In addition, the ensemble of UniBlocker and Sparkly can improve PC by up to 5% on dataset “cora”. Lastly, self-supervised dense retrieval models have outperformed BM25 as more high-quality data become available and model parameters expand. This suggests that the performance of UniBlocker will continue to improve over time, showing great promise for universal self-supervised dense blocking.
Finding 4. UniBlocker is comparable to SOTA sparse blocking methods and they complement each other to a certain extent.
@@
Furthermore, we perform efficiency analysis on scalability datasets from our proposed benchmark. To this end, we downsample the entity collections from $10^{2}$ to $10^{6}$ and compare the blocking time of different approaches. As shown in Figure 4, sparse blocking methods are efficient for small numbers of records, but the difference decreases as the number of records increases. We can see that UniBlocker has the slowest time growth curve because it generates fixed-length dense representations for records. However, for Sparkly with a 3-gram tokenizer, the number of tokens and the number of record pairs with the same tokens grow non-linearly, posing challenges for index storage and nearest neighbor search. Therefore, q-gram technology improves the effectiveness but sacrifices some scalability at larger scales.
Finding 5. Sparse blocking methods are more efficient than UniBlocker for small numbers of records, but the difference decreases as the number of records increases.
@@ 5.4. Ablation Study
We now perform ablation studies on UniBlocker to validate the importance of specific components, as depicted in Table 5.
Table 5. Average performance in ablation studies. UniBlocker “w/o OD” means pre-training with sufficient out-domain data instead of open-domain GitTables. “RE”, “LP”, “FP”, and “DD” stand for record embedding, literal paraphrasing, feature paraphrasing, and duplicate detection, respectively.
@@
UniBlocker w/o OD 46.48 89.37 18.00 35.76
UniBlocker w/o RE 48.31 89.29 19.72 33.82
UniBlocker w/o LP 49.09 89.40 20.23 34.71
UniBlocker w/o FP 48.64 89.49 19.75 35.94
UniBlocker w/o DD 48.31 89.06 19.65 36.12
UniBlocker 49.60 89.86 20.29 32.35
Effectiveness of Pre-training Data. To explore the effect of pre-training data diversity on UniBlocker, we replace the training data extracted from open-domain GitTables with out-domain tables sourced from scalability datasets in our benchmark. This helps to establish whether the pre-training data serves solely to bridge the gap between natural language and structured records, or if it also contributes to the universality of UniBlocker. Experimental results in Table 5 show that UniBlocker achieves a significant performance gain over the model trained on out-domain data, suggesting that open-domain data pre-training is crucial for its universality.
@@
Effectiveness of Data Paraphrasing. We then experiment with the effect of different data paraphrasing approaches by removing them in turn. As shown in Table 5 again, we find that both feature paraphrasing (FP) and lexical paraphrasing (LP) contribute to the performance of UniBlocker, among which LP is more pronounced because this type of paraphrasing is more refined, diverse, and stable, without the chance of introducing false positives. Moreover, due to the complexity of the real data, the application of duplicate detection (DD) achieves a 1.29% mAP improvement.
Finding 7. Pre-training on open-domain data and the techniques applied in UniBlocker are effective and necessary.
@@ 7. Conclusion
The main focus of this paper is to explore the potential of developing a universal dense blocker that can be applied across diverse scenarios and domains through large-scale self-supervised learning. To achieve this, we propose a unified contrastive pre-training framework that exploits the power of cross-domain pre-training for developing the universal dense blocker, UniBlocker. The pre-training is performed via self-supervised contrastive learning on large-scale relational tables. We also introduce a new record embedding technique that allows Transformer-based architectures to better model individual records. In addition, we propose three data paraphrasing techniques to generate positives for training. To assess the effectiveness and universality of UniBlocker, we create a new benchmark for universal blocking evaluation that covers a wide range of blocking tasks from multiple domains and scenarios. Experimental results demonstrate the effectiveness and universality of UniBlocker, holding great promise for universal dense blocking.
@@ Appendix B Benchmark Discussion
Our proposed benchmark value can be further validated by experimental results. On the one hand, Table 4 shows that existing blocking methods can effectively address the homogeneous and simple datasets that previous studies mainly focused on in Table 2. There is still room for improvement on some of our newly introduced datasets, providing new challenges and opportunities for future research. On the other hand, our proposed benchmark is the first to evaluate the general performance of different methods in different scenarios using the same parameters on all datasets, laying the foundation for future automatic, data-driven sparse blocking tuning approaches or good attribute identification for blocking. Finally, by comparing the performance differences of UniBlocker on artificially and naturally heterogeneous datasets with their corresponding homogeneous datasets, we can see that the mainstream approach of artificially constructing dirty data does not pose much of a challenge to deep models. How to better perform entity resolution directly on real highly heterogeneous data (such as “walmart-amazon_heter”) is also a future research direction.
@@ 13             $candidate\_pair\leftarrow(j,indices[j][i])$
This section presents how to use UniBlocker for blocking any number of entity record collections. It is worth mentioning that our general implementation of the nearest neighbor blocker can apply different sparse or dense vectorizers and indexers for a thorough and fair comparison of different methods. As shown in Algorithm 2, the nearest neighbor blocker operates on a list of entity collections, using a record vectorizer and a nearest neighbor indexer to identify $k$ nearest neighbors for each record. In cases with more than two collections, they are unified into a single one. The first and last collections are vectorized as the $source$ and $target$ respectively, with the latter being indexed. The algorithm iteratively searches for $k$ nearest neighbors of source vectors and stores these indices. Subsequently, candidate pairs are formulated, each constituted of a source index and its corresponding nearest neighbor index from the target. Each candidate pair is prioritized and deduplicated according to its nearest neighbor ranking, to facilitate progressive entity matching or evaluation. This sequence of candidate pair sets forms the final output of the algorithm.
Besides offering a more precise assessment for blocking, our suggested benchmark value can be further corroborated by experimental results. On one hand, Table 4 demonstrates that current sparse and dense blocking methods effectively tackle homogeneous and simple datasets primarily studied in Table 2. However, there is scope for improvement on some newly introduced datasets, which present new challenges and opportunities for future research. On the other hand, our suggested benchmark is the first to evaluate general performance across various scenarios using the same parameters for all datasets, setting the groundwork for future data-driven sparse blocking tuning approaches (Papadakis et al., 2022) or more versatile dense blocking model evaluations. Lastly, by comparing UniBlocker’s performance differences on artificially and naturally heterogeneous datasets with their homogeneous counterparts, we observe that artificial dirty data construction does not significantly challenge deep models. A future research direction includes improving entity resolution on highly heterogeneous real data (such as ”walmart-amazon_heter”).